mikewwe14:

A system of equations is shown. {2x + 2y = 17 {4x - y = 25 What is the x-value of the solution to the system of equations? Enter your answer in the box. __________

mikewwe14:

@Vocaloid

Shadow:

Mike is that you?

mikewwe14:

yes i'm on another username

Shadow:

Why's that?

mikewwe14:

it's a long long story man, i can't explain

Shadow:

I can help you in a couple min. Switching classes.

mikewwe14:

i can get vocaloid, where is she ?

Shadow:

She is busy at the moment.

dude:

Shes probably busy, Shadow is your best bet

Shadow:

\[2x + 2y = 17 \] \[4x - y = 25\] I already see a variable somewhat isolated, so lets start with finding what y is equal to in the second equation. \[4x - y = 25\] \[4x = 25 + y\] \[4x - 25 = y\] Now that we know what y is equal to, we can input this value into the second equation \[2x + 2y = 17 \] \[2x + 2(4x - 25) = 17\] \[2x + 8x - 50 = 17\] Like terms \[10x - 50 = 17\] \[10x = 67\] \[x = \frac{ 67 }{ 10 }\] \[x = 6.7\]

mikewwe14:

6.7

mikewwe14:

correct ?

mikewwe14:

can you help me with more couple questions ?

Shadow:

Sure. Just be sure to open up a new question.

mikewwe14:

you're good at math ?

Shadow:

Does it appear to be so?

mikewwe14:

just about

Shadow:

Then lets go (: